OpenStudy (jojokiw3):

Is there a method to find the outer limit of two functions when trying to find the area between two curves?

OpenStudy (freckles):

upper limit?

OpenStudy (jojokiw3):

Oop sorry. My internet sort of died. So like say i have y = 1/x and y = x^(-3/2). How do I know which one is the upper limit?

OpenStudy (jojokiw3):

Without graphing?

OpenStudy (freckles):

oh upper function?

OpenStudy (jojokiw3):

Yeah. I think.

OpenStudy (freckles):

Do you know how to find the intersections ?

OpenStudy (freckles):

also make note both functions don't exist at x=0

OpenStudy (jojokiw3):

I guess I can do 1/x = x^(-3/2)

OpenStudy (freckles):

yes

OpenStudy (freckles):

\[\frac{1}{x}=\frac{1}{x^{\frac{3}{2}} } \\ x=x^{\frac{3}{2}}\]

OpenStudy (jojokiw3):

How do you flip it?

OpenStudy (freckles):

how do i flip it?

OpenStudy (freckles):

like if 1/a=1/b then a=b as long a,b isn't 0 or you could multiply both sides by x^(3/2) and x however you like

OpenStudy (freckles):

anyways I find the zeros and were it is undefined then draw a number line and test intervals around these numbers

OpenStudy (freckles):

you should have found that the equation has a zero at x=1 and both functions are undefined at x=0 also that one with the even root doesn't exist for x<0 so we will only consider our number line for x>0

OpenStudy (freckles):

 2 intervals to check

OpenStudy (freckles):

plug in a number between 0 and 1 to see which function is larger on (0,1) plug in a number after 1 to see which function is larger on (1,inf)

OpenStudy (freckles):